OpenStudy (anonymous):

The ionization energy of hydrogen is reported to be 1.31 x 10^6 J mol^-1 of hydrogen. Compute the ionization energy for a single atom of hydrogen

OpenStudy (anonymous):

The ionization energy of hydrogen is \(1.31\times 10^6 \sf J /mol\). That means the ionization energy for a single H atom is \(\large \sf 1.31\times 10^6 J/mol \times \frac {1~mole~H}{6.022\times 10^{23}~H~atoms}=...J/atom\)